A company with 2 million shares of common stock currently outstanding is planning to sell 500,000 new shares to its existing shareholders through a rights issue. The current market price of a share is $65, and the subscription price is $55. If the stock is selling rights-on, calculate the number of rights needed to purchase one of the new shares of common stock and the value of each right.

a. Calculate the number of rights needed to buy one share
b. Calculate the value of a right. $
Answer #1

a

number of rights needed to buy one share = current shares/new shares = 2000000/500000 = 4

b

value of a right = (current market price-subscription price)/number of rights needed to buy one share

=(65-55)/4 = 2.5